OpenStudy (blackalmaz):

Help me prove this!Will medal! Prove using the definition of the limit: lim as x approaches -2 of (x^2 + 3x - 1) = -3

OpenStudy (blackalmaz):

\[\lim_{x \rightarrow -2} (x^2 + 3x -1) = -3\]

OpenStudy (zzr0ck3r):

this is a continuous function on the point in question. i.e. plug it in

OpenStudy (blackalmaz):

Yeah, I know the limit's value but I need to prove it using the definition of the limit. Definition of limit: Let f(x) be a function defined on an interval that contains x=a, except possibly at x=a. Then we say that, \[\lim_{x \rightarrow a}f(x) = L\] if for every number epsilon > 0, there is some number delta > 0, such that \[\left| f(x) - L \right| < \epsilon , whenever 0 < \left| x- a \right| < \delta\]

OpenStudy (agent0smith):

Ew. Haven't done these in a while. \(\large |(x^2+3x-1) - (-3)| < \epsilon\) wherever \(\large 0 < \left| x- (-2) \right| < \delta\) simplify and factor \(\large |x^2+3x+2| < \epsilon\) wherever \(\large 0 < \left| x+2 \right| < \delta\) \(\large |x+2| |x+1| < \epsilon\) wherever \(\large 0 < \left| x+2 \right| < \delta\)

OpenStudy (blackalmaz):

I know about that, but you need to get rid of x+2 in order to prove it.That's where I'm having trouble.But thanks.

OpenStudy (blackalmaz):

I meant x+1.

OpenStudy (irishboy123):

its continuous you "prove it" by plugging it in

OpenStudy (blackalmaz):

No, in this case "proving" means using the definition of the limit, to riguriously state that. Simply saying that that's true doesn't prove it.

OpenStudy (agent0smith):

Example 3: http://tutorial.math.lamar.edu/Classes/CalcI/DefnOfLimit.aspx

OpenStudy (irishboy123):

you can plug in \(x = (- 2 + h)_{h \to 0} \) if you like with \(0 < |h| << 0\) but not here

OpenStudy (blackalmaz):

Lol Paul's online math notes are exactly where I am learning from. I don't quite get their explanation. Why do we set x+5 smaller than some number K?

OpenStudy (agent0smith):

\[\large k |x+1| < \epsilon\]so that you can do it, like all the other examples, and write it as \[\large |x+1| < \frac{ \epsilon }{ k } \]

OpenStudy (blackalmaz):

Yeah, but here you substituting K for the x+5, isn't K bigger than x+5?

OpenStudy (agent0smith):

Oops, should be x+1

OpenStudy (blackalmaz):

Oh, I get it now. I have another question, why do you pick the minimum of the 1, epsilon/10?

OpenStudy (agent0smith):

We're using \(\large |x+1|<k\) so that \[\large |x+2||x+1|<k|x+2|\]

OpenStudy (agent0smith):

I think Paul's notes explain that well.

OpenStudy (blackalmaz):

What does guaranteeing that delta is smaller than or equal to epsilon/10 do? Everything doesn't quite piece together in my head.

zepdrix (zepdrix):

So like